# Enhancing Robustness in Language-Driven Robotics: A Modular Approach to Failure Reduction

### v1

@@ IV Experiments
We test our framework in two different settings: a service robot setting and an industrial manipulation setting. We refer to https://tinyurl.com/4h2ysybd for more details on our experiments and our full code base.
@@ IV-A Service setting
We consider a mobile manipulator robot with one arm (for example a TiaGO[33] robot), helping a user in a home. We use a corpus of ten pick-and-place tasks which are typical of what a user might require from a service robot:
  • ”Bring me something to water my plant with. There is a water glass on the kitchen table and I am at the coffee table.”
  • ”Bring pills and a water glass to my desk. The items are on the bathroom counter. The robot is in the bathroom.”
  • ”I’m hungry. There is a peach on the coffee table and I am at my desk.”
  • ”Bring me a drink. There is a water glass on the kitchen table. The robot is at the user desk and the gripper is empty.”
  • ”Bring the mouse to the table and put the mug on the kitchen counter. The mouse and the mug are on on the desk.”
  • ”Put the pliers and screwdriver on the shelf and the bottle in the trash. The objects are currently on the desk.”
  • ”I am done eating. Please clear the table. There are a wine glass and a plate on the table.”
  • ”Bring me something to clean this stain on the table. There are soap and a sponge on the kitchen counter.”
  • ”Put away the pills in the bathroom cabinet. The pills are on the bathroom counter.”
  • ”Put the pie in the trash can.”
The experiments are run using the VirtualHome simulator [34, 35]. In VirtualHome, the robot is represented by a humanoid agent. For realism, we add a $10\%$ chance of failure in the grasp primitive. Experiments are run until the plan runs out of steps or the amount of steps becomes double the amount of steps in the plan (in which case we consider the run to have timed out and the task to be failed).
For each task, we also use GPT-4o to generate five alternative task statements. We instruct GPT to introduce variety in the statements, while ensuring that both the task and the information about the initial of the objects is included. We solve these tasks using the full architecture to verify its suitability to non-expert users.
@@ IV-B Industrial setting
Manipulation tasks: We consider a robot arm carrying out tasks inspired by the robothon task board [36]:
@@
In this setting, the manipulation tasks can only be carried out if the object is properly grasped before attempting to fulfil the task. For the charger task, the charger must be grasped by the plug part, while for the probe task, the probe must be grasped by the handle. We allow the agent to specify the part of the object on which the grasp is carried out by adding a parameter to the grasp primitive.
User-oriented tasks: We ask ten participants to ask the robot to solve the task board. We describe the task requirements but provide no details on the robot’s skills. We also design a baseline task statement for comparison. For these experiments, the primitives represent the skills required by the task board: press a blue button, plug the probe cable and open the trapdoor, where the last two skills were obtained from the manipulation tasks above.
@@ IV-C Ablation study modalities
  • CaP: Inspired by Code-as-Policies[7], the LLM only has access to the task and to the primitives and must output the full sequence of skills. We use GPT-4o for this version.
  • Plan: There is no expected outcomes module, and the execution module’s output is directly executed on the robot, without verification or feedback. We use LLaMA3.1 for the planner and exection module.
  • EO: Expected outcomes: As above, there is no feedback. The expected outcomes module is active and the expected outcomes are given to the execution module when it is initialized and at each time step. All modules use LLaMA3.1.
  • FB: The execution module receives feedback, but the expected outcomes module is not active. As for the plan-only experiments, the current step is made explicit. All modules use LLaMA3.1.
@@
V Results
In this section, we show that our architecture significantly outperforms the baselines on our task corpora.
@@ V-A Service robot results
We run each experiment five times, comparing the success rates of each variant of the architecture. A mean, std plot of the success rates can be found in Figure 6.
Over all tasks, only our full architecture reaches a success rate of $100\%$. The components of our architecture improve the success rate over the plan baseline but still fail in most cases. Interestingly, the expected outcomes module drastically reduces the rate of syntax errors in the execution module’s output: the proportion of steps that failed due to incorrect syntax was $0.10$ for the full architecture, $0.51$ for the FB architecture, $0.16$ for the EO architecture and $0.57$ for the plan architecture. This significant increase in code quality significantly contributes to the full architecture’s success rate. Simulation snapshots taken from a successful execution of the task ”Bring me something to water my plant with. There is a water glass on the kitchen table and I am at the coffee table.”, using our full architecture, are shown in Figure 5.
Figure 5: Successful execution of the ”Bring me something to water my plant with. There is a water glass on the kitchen table and I am at the coffee table.” task. From left to right, top to bottom: (i) the service robot approaches the kitchen table; (ii) the robot grasps the glass of water; (iii) the robot moves to the coffee table, (iv) the task ends with the robot putting down the glass. In VirtualHome, the agent is represented by a humanoid figure.
Despite GPT-4o typically performing better when generating robotic code, the code-as-policy approach obtains a lower success rate than our architecture. The two main causes for this are a lack of alignment between the LLM and the actual task and the open-loop execution failing to recover from grasp failure.
The success rates with our full architecture drops to $72\%$ with the GPT-4o generated task statements. Failures arise from tasks stated in a way similar to a plan, leading the planner to mirror the structure of the task statement and outputting an unfeasible plan. Errors also happen when the task is stated with a vocabulary that is very different from the primitives. In such cases, the execution module fails to use the correct skill to carry out the plan. We leave the exploration of tradeoffs between open-vocabulary planning and ensuring a degree of closeness between the plan and the primitives or future work.
Figure 6: Mean success rates and standard deviations (black lines) over the service robots task corpus. Bars in blue denote architectures using LLaMA3.1, the red bar th architecture using GPT-4o. Note that the full architecture (right) has no standard deviation, as it recovers from grasp failure.
@@ V-B Industrial robot results
Manipulation tasks: Both manipulation tasks are successfully completed by both versions of our architecture that use the expected outcomes module and by the baseline using GPT-4o. In experiments without the expected outcomes module, LLaMA3.1 uses the generic grasp skill or erroneous parametrizations of it, resulting in failure of the subsequent skills due to unsuited grasps (see Figure 7 for an example of successful execution). This demonstrates the expected outcome module’s ability to correctly specify the subgoals.
@@
Figure 7: Successful execution of the ”Plug in the charger” task. From left to right, top to bottom: (i) the gripper approaches the charger’s plug; (ii) the charger is grasped by the gripper; (iii) the arm begins plugging the charger, (iv) the task ends with the charger plugged
Taskboard solution: Our pipeline succeeds in solving the taskboard when we use the expert-designed baseline prompt. Out of $10$ participants, $9$ prompts lead to successful taskboard solutions. The failure stems from the participant’s instructions being overly complex and detailed, leading the planner to output a plan where some steps aren’t feasible with the robot’s skills (for example, the prompt instructs the robot to ’Check that the door is opened’, which isn’t aligned with the robot’s skills).
VI Discussion
Our experimental corpus shows the improvement in performance of our architecture compared to both the baseline and ablated versions of the pipeline. Our service robot experiments demonstrate that feedback mechanisms allow the agent to recover from misalignments between the planner and the primitives. Intriguingly, we also find that explicitly characterizing the execution and outcome of each plan step leads to the execution module outputting higher quality code, with a significant decrease in syntax errors. In our industrial robot experiment corpus, we show how the expected outcomes module can help properly define subgoals: by adding detail to the execution module’s input, information missing from the plan can be recovered. Finally, by obtaining alternate task statements (through GPT-4o in the service domain and through participants in the industrial domain), we show that our pipeline retains high success rates when the task is specified by a non expert with limited knowledge. This is an encouraging insight into the potential of LLMs for language-driven robotics.
VII Concluding remarks and future work
@@
In this paper, we introduced a formalism for the problem of planning and executing robotic tasks with language models as the problem of following a set of subgoals through a chain of options in a goal-conditioned POMDP. We showed that execution errors of such a pipeline can be classified within this formalism by remarking that they amount to discrepencies between subgoals and some ground-truth subgoals. Then, we proposed an architecture with technical components designed to prevent such errors. Namely, we propose that adding a second planning component designed to output a detailed plan and implementing some feedback mechanism leads to a sharp increase in performance. Through experiments in a service robot setting, we show that our architecture leads to significantly more robust execution for pick-and-place tasks, and experiments in an industrial robot setting show that our pipeline also succeeds at manipulation tasks. Our hardware validation shows that our architecture can be applied in real time using a standard computer and an openly accessible language model, providing encouraging insights into the applicability of smaller language models for robotics.
@@
Future research directions include using our architecture to facilitate the acquisition of new skills and testing our pipeline as part of a cognitive architecture for a service robot.

### v2

@@ IV Experiments
This section describes our experimental protocol, in two robot settings: service and industrial.
@@ IV-A Service setting
We consider a one-armed mobile manipulator robot helping a user in a home. The robot is used to carry out pick-and-place actions as might be required of a robotic home assistant. We randomly generate a set of $50$ pick-and-place tasks, each involving a random item, original item location and target item location. For each task, we programatically generate the instruction, randomly sampling among $3$ templates, and we use GPT-4o to generate two alternate task instructions. We instruct GPT to introduce variety in the statements, while ensuring that both the task and the information about the initial position of the objects are included. $10$ randomly selected instructions can be found in the appendix. We run an ablation study in order to verify the usefulness of our modules, as detailed in Section IV-C.
The experiments are run using a symbolic simulator inspired by VirtualHome [36, 37], and we use VirtualHome to render videos of our experiments (see snapshots in Figure 4). In VirtualHome, the robot is represented by a humanoid agent. We implement a middleware between our architecture and the simulator, where the motion primitives and corresponding feedback mechanisms are implemented. We add a $10\%$ chance of failure in the grasp primitive. Experiments are run until the task is completed or the amount of interactions between the execution module and the environment becomes double the amount of steps in the plan (in which case the run has timed out and the task is failed).
@@ IV-B Industrial setting
We consider a robot arm carrying out manipulation tasks inspired by the robothon task board [38]. This includes $2$ manipulation tasks and the full task board solution (press the blue button, plug in the probe cable, press the red button and open the trapdoor). The experiments in this setting were validated on a Franka arm, see Figure 5 and the attached video (available on our git).
Manipulation tasks: We consider the following manipulation tasks
@@
These tasks can only be carried out if the object is properly grasped before attempting to fulfil the task. For the charger task, the charger must be grasped by the plug part, while for the probe task, the probe must be grasped by the handle. We allow the agent to specify the part of the object on which the grasp is carried out by adding a parameter to the grasp primitive. We attempt the manipulation tasks using the Full, FB, EO and Plan versions of our architecture for these tasks.
Task board: We ask ten participants to ask the robot to solve the task board. We describe the task requirements but provide no details on the robot’s skills. For these experiments, the primitives available to the robot represent the manipulation skills required by the task board: press a button, plug the probe cable and open the trapdoor, where the cable plugging skill was obtained from the manipulation task above, and we use our full architecture.
@@ IV-C Ablation study modalities
  • CaP: Inspired by Code-as-Policies[7], the LLM receives the instruction and the primitives and must output the full sequence of skills. We use GPT-4o for this approach.
  • Plan: There is no expected outcomes module, and the execution module’s output is directly executed on the robot, without verification or feedback. We use LLaMA3.1 for the planner and execution module.
  • Expected outcomes (EO): As above, there is no feedback. The expected outcomes module is active and the expected outcomes are given as input to the execution module. All modules use LLaMA3.1.
  • FB: The execution module receives feedback, but the expected outcomes module is not active. Both modules use LLaMA3.1.
@@
  • DS full: Additionally, we implement a variation of the Full architecture, where we use the Deepseek-r1:7b model instead of LLama. Aside from a few modifications to the parsing mechanisms to accommodate Deepseek’s output format, the code is the same as in the Full experiments.
V Results and discussion
In this section, we show that our architecture significantly outperforms ablated versions, even surpassing the GPT-based code-as-policies baseline. We also demonstrate its suitability to real robot use through a hardware experiment. By studying the amount of compute (average tokens and execution module inferences) required for the service tasks, we show how our modules result in faster interaction, facilitating data generation.
@@ V-A Service robot results
We run the experiment corpus three times, using the different instruction sets. A mean, std summary of the success rates of each architecture variant can be found in Table I.
Over all tasks, our full architecture strongly outperforms its ablated versions. Individually, the components of our architecture each improve the success rate over the plan version. Interestingly, we notice that by precisely characterizing each subgoal, the expected outcomes module increases the code quality of the execution module’s output, leading to a sharp increase in success rate. Simulation snapshots taken from a successful execution of the task ”Move the Water Glass to the Coffee table. It is currently on the Kitchen table.”, using our full architecture, are shown in Figure 4.
Figure 4: Successful execution of the ”Move the Water Glass to the Coffee table. It is currently on the Kitchen table.” task. From left to right, top to bottom: (i) the service robot approaches the kitchen table; (ii) the robot grasps the glass of water; (iii) the robot moves to the coffee table, (iv) the task ends with the robot putting down the glass.
Despite GPT-4o’ typical performance at code generation tasks, the code-as-policy approach obtains a lower success rate than our architecture. The two main causes for this are a lack of grounding between the LLM and the environment and the open-loop execution failing to recover from grasp failure. We also note that the Deepseek-based architecture almost never succeeds. We conjecture that this is due to the model being specialized in one-shot reasoning and problem solving, making it less suitable for long-horizon interactions.
Architecture Success rate Execution calls Tokens output
CaP 92.67 $\pm$ 3.40 1 $\pm$ 0 88.13 $\pm$ 0.98
Plan 10.67 $\pm$ 3.40 4.11 $\pm$ 0.02 386.67 $\pm$ 15.11
EO 87.33 $\pm$ 0.94 4.11 $\pm$ 0.02 283.20 $\pm$ 11.46
FB 67.33 $\pm$ 6.18 6.11 $\pm$ 0.12 547.40 $\pm$ 73.71
Full 96.00 $\pm$ 2.83 5.19 $\pm$ 0.47 320.00 $\pm$ 30.80
Full DS 17.33 $\pm$ 0.94 6.96 $\pm$ 0.22 3627.67 $\pm$ 241.17
Table I: Percentage success rate (mean $\pm$ std), average number of calls to the execution module and average number of tokens output by the execution module for each of the architectures in our benchmark. Our full architecture outperforms ablated versions, and surpasses the GPT-based code-as-policies method. Both the expected outcomes module and feedback mechanism increase success rate, and the expected outcomes module reduces the amount of execution module calls and the number of tokens output by the model, resulting in faster execution.
@@ V-B Industrial robot results
Our full pipeline succeeds in solving the task board with each of the $10$ participants instructions, and in carrying out both manipulation tasks. When using the Plan and FB architectures, the success rate in the manipulation tasks drops to $0.5$, due to the execution module failing to properly specify which part of the object should be grasped. This highlights how the expected outcomes module can be used to improve subgoal characterization (in this case by specifying grasp targets on the objects), leading to more robust execution.
These hardware experiments (see Figure 5 and the video attachment to this paper) show that our architecture is suitable for real-time use on a real robot.
@@
Figure 5: Successful execution of the ”Plug in the charger” task. From left to right, top to bottom: (i) the gripper approaches the charger’s plug; (ii) the charger is grasped by the gripper; (iii) the arm begins plugging the charger, (iv) the task ends with the charger plugged.
V-C Computation and inference time
In our service experiments, when using our full architecture, our execution module outputs on average $3.64$ times more tokens than the GPT-based architecture (see Table I). It is estimated GPT-4o is a mixture of $110$ billion parameter experts (MoE), as opposed to the $8$ billion of LLama3.1, which is $13.75$ times smaller. As the amount of floating-point operations is approximately linear with the model (or expert) size and the number of output tokens, it results that our method needs about 3.78 times less compute operations per task compared to a GPT-based architecture. Further, our execution module inference time is on average 2.05 seconds on a standard ($8$GB VRAM) GPU, showing the suitability of our approach for generating large datasets and for online robotic interaction.
VI Concluding remarks and future work
@@
In this paper, we formalized the problem of planning and executing robotic tasks with large language models as the problem of following a set of subgoals through a chain of options in a goal-conditioned POMDP. We showed that execution errors of such a pipeline can be classified within this formalism as missing or mischaracterized subgoals preventing option chaining. Then, we proposed an architecture with technical components designed to prevent such errors. Namely, we added a second planning component designed to explicitly characterize the outcome of each plan step and implemented some feedback mechanisms. Through experiments in a service robot setting, we showed that our architecture achieves robust execution, while incurring low computational costs (3.78 times lower than a GPT-based baseline), for pick-and-place tasks, and experiments in an industrial robot setting show that our pipeline also succeeds at manipulation tasks. Our hardware validation shows that our architecture can be applied in real time using a standard computer and an openly accessible language model, providing encouraging insights into the applicability of smaller language models for robotics.
@@
Future research directions include testing our architecture as part of a robotic cognitive architecture and fine-tuning models for robotic planning with data obtained using our pipeline.
@@ Acknowledgements
Appendix A Prompts
We report the prompt templates used for our service experiments. The prompts used in the industrial setting are the same up to minor adaptations, and can be found on our github. The colored text segments are programatically added to the prompts:
  • Task: Task instruction
  • Plan: Plan generated by the planning module, as a Python-style list of strings
  • EOs: Expected outcomes generated by the EO module, as a Python-style dictionary of strings
  • Skills: Python-style headers for the functions containing the motion primitives
  • Environment: Name of the objects and locations in the environment (in Python string format), and location of the objects and robot at the beginning of the trial.
Planner module prompt:
You are in charge of a mobile robot with an arm ending in a gripper. Your task is the following: Task
Please output an plan, composed of simple actions, to carry out this task. Remember that the robot should always move to a location before interacting with objects in this location, unless it is already there. However, you can assume that simple actions (such as grasping or putting down objects) automatically move the arm to the correct position.
Please only output the plan as a tuple of strings, where each step is a string, without any other text.
Expected outcomes module prompt:
You are in charge of executing the following task: Task. The plan consists of the following steps: Plan Each of the steps of the plan will be executed with a mobile robot equipped with an arm ending in a gripper. For each step of the plan, I need you to give the expected outcome of the actions involved in the step, in physical and visual terms.
This should consist of one or two short, simple sentences that are a more complete and detailed description of the step’s outcome. The sentences should describe the final state of the robot, for example if it should be at a location, have grasped an object (and what part of the object, if relevant for the task), or where an object should be put down. You can add some information if the plan is too concise. Here are some examples, with the plan step first and the expected outcome after:
- Put bottle on shelf: The bottle should be on the shelf.
- Grasp the mug: The mug should be in the robot’s gripper.
- Grasp the knife: The knife blade should be in the robot’s gripper.
For each step of the plan, please briefly describe the expected outcome as shown above. Please try to be concise and focus on the most relevant information. Please fill out the following python dictionary with the expected outcomes: dictionary template. Only output the dictionary and no other text.
Execution module prompt:
Context:
Your are now in charge of a mobile robot equipped with one arm with a parallel gripper. You will be given a high-level task that you will need to fulfill using this robot, and the corresponding plan, which is a series of simpler steps. You will need to carry out the task step by step by interacting with the system using some code primitives. At each step the plan will be updated and you will receive feedback. The skills are python functions, which allow you to perceive and act on your environment.
Skills:
Here are the functions and skills, with examples of the syntax:
skills
The task and the plan: You are in charge of executing the following task: Task. The plan is: Plan
Here are the expected outcomes of each step in the plan, which you can use as a guide:
EOs
Environment
What I need you to do:
Please define a function do(), which will contain mostly action primitives to solve the steps of the plan one by one. Please output python code, enclosed between the tags <code> and </code>. Please only use the functions I defined above and ensure the locations and objects that you pass as arguments are correct.
Appendix B Tasks
Here are $10$ randomly-selected sample task instructions from the service robot corpus:
  • Move the Water Glass to the Coffee table. It is currently on the Kitchen table.
  • Move the Pills from the Desk to the Kitchen counter
  • Move the Fork to the Kitchen table. It is on the Coffee table.
  • Put the Mouse on the Table. The Mouse is on the Desk.
  • Transfer the Knife from the Desk to the Kitchen table.
  • Set the Screwdriver down on the Coffee table. It is now on the Desk.
  • Place the Plate onto the Kitchen counter. It is on the Coffee table.
  • Retrieve the Knife from the Coffee table and place it on the Desk.
  • Take the Cupcake and put it on the Desk. The Cupcake is on the Table.
  • Move the Fork onto the Table. It is on the Desk.